OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = 9 –3x + 2y = 9

OpenStudy (anonymous):

3x-4y=9 -3x+2=9 Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so add them, any idea how?

OpenStudy (anonymous):

second one is -3x+2y=9*

OpenStudy (anonymous):

I'm so confused. Idk how to do any of this. I'm so terrible at math.